# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

### v1

@@ 2.2 Bayesian Inference for Data-driven Discovery
With the Regularized Horseshoe prior, most identified coefficients are shrunk aggressively toward zero, which effectively removes irrelevant or redundant basis functions. On the other hand, significant coefficients escape shrinkage due to the slab, which preserves their magnitude. The inverse-Gamma prior on $c^{2}$ ensures that the slab variance remains bounded, which avoids failed coefficient identification in MCMC sampling. We will elaborate on how we employ gradient-based MCMC methods in both identifying system coefficients and provide robust UQ within limited data in the next section.
3 Langevin-Assisted Active Physical Discovery
In this work, we introduce LAPD, which incorporates Langevin MCMC methods into the discovery of physical laws in a Bayesian way via a sequential threshold method. LAPD utilizes the stochastic gradient in the situation of sufficient data collection to improve efficiency and can be exploited for AL in the face of data scarcity (i.e., the time derivative calculation or collection is expensive). We start with the data collected $\mathbf{U}\in\mathcal{R}^{N\times d}$ from the ODE or PDE systems as described before, with the task of identifying the system in terms of (1) and determining the posterior distribution of the model coefficient $\mathbf{\Xi}\in\mathbb{R}^{k\times d}$, where $k$ is the number of candidates in each row of the library $\mathbf{\Theta}(\mathbf{U})$.
3.1 Uncertainty Quantification with Langevin MCMC
$d\mathbf{\Xi}_{t}=-\nabla U(\mathbf{\Xi}_{t})\,dt+\sqrt{2\tau}\,d\mathbf{W}_{t},$ (4)
where $U(\cdot)$ represents the energy function, $\{\mathbf{W}_{t}\mid t\geq 0\}$ denotes the standard Brownian motion on $\mathbb{R}^{m\times d}$, and $\tau>0$ is the temperature parameter. Under mild conditions on $U$, it is well-established that the diffusion process described by (4) admits a unique strong solution $\{\mathbf{\Xi}_{t},t\geq 0\}$, which is a Markov process. Moreover, as $t\to\infty$, the distribution of $\mathbf{\Xi}_{t}$ converges to the invariant distribution $\pi_{\tau}$, characterized by the density $\pi_{\tau}(\mathbf{\Xi})\propto\exp(-U(\mathbf{\Xi})/\tau)$.
$\tilde{\mathbf{\Xi}}_{k+1}=\tilde{\mathbf{\Xi}}_{k}-\eta_{k}\nabla U(\tilde{\mathbf{\Xi}}_{k})+\sqrt{2\eta_{k}\tau}\,\zeta_{k},$ (5)
where $\zeta_{k}\sim\mathcal{N}(0,I_{md\times md})$ represents independent Gaussian noise at each iteration, and $\eta_{k}$ is the step size at iteration $k$. To enable scalable and efficient sampling while maintaining convergence to the target distribution under appropriate conditions, we further consider its stochastic gradient version for efficient large-scale data applications. Specifically, the energy function is approximated with a mini-batch of data $\mathbf{B}$ uniformly subsampled from the given data (Welling and Teh,, 2011):
$\tilde{L}(\tilde{\mathbf{\Xi}})=-\log p(\tilde{\mathbf{\Xi}})-\frac{N}{|\mathbf{B}|}\sum_{\mathbf{x}_{i}\in\mathbf{B}}\log P(\mathbf{x}_{i}\mid\tilde{\mathbf{\Xi}}),$ (6)
Given the problem setup (2), the gradient $\nabla\tilde{L}(\tilde{\mathbf{\Xi}})$ is calculated as
$\nabla\tilde{L}(\tilde{\mathbf{\Xi}})=-\nabla\log p(\tilde{\mathbf{\Xi}})-\frac{N}{|\mathbf{B}|}(\mathbf{\Theta}(\mathbf{B})^{\intercal}(\dot{\mathbf{B}}-\mathbf{\Theta}(\mathbf{B})\tilde{\mathbf{\Xi}})).$ (7)
For an appropriately chosen step size schedule $\eta_{k}$, it has been established that the Langevin MCMC converges to the target stationary distribution (Durmus and Moulines,, 2018). Despite the theoretical guarantees, the existing non-asymptotic convergence results can be difficult to interpret and apply in practice, particularly when step size schedules must balance convergence speed with accuracy. To address these challenges, several methods, such as underdamped Langevin MCMC (Cheng et al.,, 2018), Hamiltonian Monte Carlo (Neal et al.,, 2011), cyclical SGMCMC (Zhang et al.,, 2019), reSGLD (Chen et al.,, 2018), etc., have been developed as extensions of the Langevin MCMC framework. These methods aim to enhance computational efficiency and scalability, especially in high-dimensional settings.
We consider reSGLD to balance exploration and exploitation when identifying system coefficients, which offers a blend of efficiency and robustness in traversing different energy levels to address large-scale non-convex sampling problems (Deng et al.,, 2020; Zheng et al.,, 2024). It simulates a high-temperature chain for exploration and a low-temperature chain for exploitation. The sampling process is shown as follows:
$\tilde{\mathbf{\Xi}}_{k+1}^{(1)}$ $=\tilde{\mathbf{\Xi}}_{k}^{(1)}-\eta_{k}\nabla\tilde{L}(\tilde{\mathbf{\Xi}}_{k}^{(1)})+\sqrt{2\eta_{k}\tau_{1}}{\zeta}_{k}^{(1)}$ (8)
$\tilde{\mathbf{\Xi}}_{k+1}^{(2)}$ $=\tilde{\mathbf{\Xi}}_{k}^{(2)}-\eta_{k}\nabla\tilde{L}(\tilde{\mathbf{\Xi}}_{k}^{(2)})+\sqrt{2\eta_{k}\tau_{2}}{\zeta}_{k}^{(2)},$
where $\tilde{\mathbf{\Xi}}_{k}^{(1)}$ and $\tilde{\mathbf{\Xi}}_{k}^{(2)}$ denote the sampling results of two chains with temperatures $\tau_{2}>\tau_{1}$. Furthermore, we swap the Markov chains in (8) with the corrected swapping rate for mini-batch setting:
$\tilde{S}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)},\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}=e^{\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\left(\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)}\big{)}-\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}-\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\frac{\tilde{\sigma}^{2}}{C}\right)}$ (9)
where $\tilde{\sigma}^{2}$ approximates the variance of $\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)}\big{)}-\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}$ and $C$ acts as an adjustment to balance acceleration and bias. The algorithm proceeds iteratively until a specified number of iterations is reached. Generate uniform $u\in[0,1]$, if $u<\tilde{S}$, then perform the swapping between two chains. The parameter sets $\{\tilde{\mathbf{\Xi}}_{k}^{(1)}\}_{k=1}^{K+1}$ are produced as outputs for analytical purposes.
3.2 Bayesian Physical Discovery with Limited Data
The integration of AL into data driven discovery of dynamical systems attempts to address two fundamental challenges inherent to learning parsimonious dynamical models from limited and costly data. First, accurate estimation of state derivatives $\mathbf{\dot{U}}$, which is critical for regression-based identification of (1), becomes unreliable when the data are irregularly sampled or sparsely distributed. Traditional numerical differentiation (e.g., finite differences) amplifies noise and errors under these conditions, particularly when time intervals are large or nonuniform. Second, dynamical systems often exhibit heterogeneous informativity across their state space: sparse governing terms dominate in specific regimes (e.g., near bifurcations or transient states), while other regions contribute minimally to identifying the dynamical systems. Passive or uniform sampling risks undersampling these critical regions, leading to biased or even incorrect models.
AL directly targets these issues by prioritizing data acquisition, where the uncertainty in the learned dynamics is maximized. By tackling (2) within a Bayesian framework, posterior uncertainties over the coefficients $\mathbf{\Xi}$ in (2) can be quantified. Regions of high predictive entropy or variance in $\mathbf{\dot{U}}$ correspond to states in which the model cannot confidently distinguish between candidate terms in the library $\mathbf{\Theta}(\mathbf{U})$. Actively querying derivatives in these states maximizes the information gain on $\mathbf{\Xi}$, which efficiently resolves ambiguities in the sparsity pattern while minimizing redundant sampling (Riis et al.,, 2023; Pickering et al.,, 2022; Gramacy and Lee,, 2009; Fasel et al.,, 2022).
To formulate AL, we first denote $\mathcal{D}=\{\mathbf{U},\dot{\mathbf{U}}\}$. With the information of the posterior $p(\bm{\Xi}|\mathcal{D})$,
@@ 3.2 Bayesian Physical Discovery with Limited Data
During implementation, we estimate $\mathbb{E}_{p(\boldsymbol{\bm{\Xi}}\mid\mathcal{D})}\left[\sigma_{\boldsymbol{\bm{\Xi}}}^{2}(\dot{\mathbf{u}})\mid\boldsymbol{\bm{\Xi}}\right]$ with the predictive variance $\sigma^{2}(\dot{\mathbf{u}})$. The higher the variance, the more uncertainty we have for the sample.
@@
We sample $P$ groups of the posterior distribution of model parameters and calculate $\dot{\mathbf{u}}^{pred}_{i}$ for each model $i$. The predictive variance is calculated as
$\sigma^{2}({\dot{\mathbf{u}}})=\frac{1}{P}\sum_{i=1}^{P}(\dot{\mathbf{u}}^{pred}_{i}-\bar{\dot{\mathbf{u}}}^{pred})^{2},$
leading to the selection of potential design points with high uncertainty. Through experiments, it is observed that using the uncertainty acquisition function alone may cause the selected points to cluster in different rounds. While those points are informative, we are missing information from other regions, which causes a rather biased model favoring and overfitting one region of the data.
To solve this problem, we combine the above-mentioned uncertainty acquisition function with a space-filling design to spread the design points. Here we adopt the maximin distance criterion (Loeppky et al.,, 2010; Yu and Kim,, 2010), which selects the next design points that maximize the minimum distance to the current design points. Here, the maximin criterion is adjusted by the density of the data.
$d(\mathbf{\mathbf{\Theta}(\mathbf{u})})=\text{min}||\mathbf{\Theta}(\mathbf{u})-\mathbf{\Theta}(\mathbf{u}_{i})||_{2}\cdot\text{density}(\mathbf{u})^{\lambda},\text{for all}~{}\mathbf{u}_{i}\in\mathcal{M},$ (10)
where $\lambda\geq 0$ is a tunable hyperparameter that controls how much density penalizes on the distance. when $\lambda=0$, the criterion returns to the maximin criterion with no adjustment by density. Density-based adjustment has been adopted in AL research for both classification and regression tasks to achieve different goals, such as increasing the representativeness of the data (Wang et al.,, 2021), mitigating the sensitivity of the Euclidean distance to the scale and distribution of the data (Donmez and Carbonell,, 2007) and preventing outliers added to the training data (Zhu et al.,, 2008). The density measure includes KDE, cosine distance-based measurement, and K-Nearest-Neighbor-based density measure. We opt for the K-Nearest-Neighbor-based density measure over KDE or cosine similarity because KNN offers computational simplicity and directly captures local data density without making strong assumptions about the underlying data distribution, and unlike cosine similarity, it preserves both directional and magnitude information that is essential for identifying truly representative samples in our feature space. The K-Nearest-Neighbor-based density is calculated as the inverse of one point’s average distance to its K nearest neighbor.Our goal here is to penalize on data whose density is low, since in this situation, there is a higher chance that it holds high maximin distance with data already selected, especially when the data itself is also of high magnitude. In the sense that without density adjustment, data of low density and high magnitude are preferred, while data with high density and high magnitude are disregarded. This prevents the criterion from achieving the goal of space-filling and potentially favoring the same group of data as the uncertainty criterion. When the data distribution is quite uniform or the magnitude does not vary too much, there is no need to do a density adjustment.
Two criterion are standardized into $[0,1]$ to get $\tilde{\sigma^{2}}({\dot{\mathbf{u}}})$ and $\tilde{d}(\mathbf{\mathbf{\Theta}(\mathbf{u})})$ and the new acquisition function is defined by weighted sum, with $\alpha$ being the tunable weight,
$C(\mathbf{u})=\alpha\,\tilde{\sigma^{2}}({\dot{\mathbf{u}}})+(1-\alpha)\,\tilde{d}(\mathbf{\mathbf{\Theta}(\mathbf{u})}).$ (11)
@@
Input: Pool of potential design points $\mathcal{D}$; number of initial points $n$; batch size of selected points per round $m$, maximum allowed sample size $N_{max}$, tolerance of convergernce Tol.
@@
Randomly select $n$ points, $\mathbf{u}_{1},\dots,\mathbf{u}_{n}\in\mathbf{U}$ with derivatives $\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{n}$ available;
Set of selected points =$\mathcal{M}=\{(\mathbf{u}_{1},\dot{\mathbf{u}}_{1}),\dots,(\mathbf{u}_{n},\dot{\mathbf{u}}_{n})\}$, $\mathcal{D}\leftarrow\mathcal{D}/\mathcal{M}$;
@@
       Use $\mathcal{M}$ points to construct the library $\mathbf{\Theta}$ and perform LAPD to estimate the posterior distribution $p(\boldsymbol{\bm{\Xi}}\mid\mathcal{M})$;
@@
Algorithm 1 Active Learning in LAPD: Iteratively improve the model by selecting design points that maximize the acquisition function, updating the posterior, and retraining the model until convergence.
Intuitively, the synergy arises from AL’s ability to exploit the structured sparsity of dynamical systems. Unlike generic regression tasks, SINDy can identify a small subset of relevant terms from a potentially large library. Correlated or redundant terms create degeneracies that passive data may fail to disentangle. AL breaks these degeneracies by selecting states in which the predictions of competing sparse models diverge most significantly. To make it clearer, in a system with cubic nonlinearity terms, this strategy will favor sampling in high-energy regions where the cubic terms dominate rather than in low-energy regions where the linear-term approximations suffice. It not only reduces the number of required $\mathbf{\dot{U}}$ evaluations but also enhances robustness to noise and model misspecification, which ensures that the identified dynamics are both parsimonious and physically consistent.
3.3 The Proposed Algorithms

### v2

@@ 3.1 Uncertainty Quantification with Langevin MCMC
$),\nabla\tilde{L}(\tilde{\bm{\xi}})=-\nabla\log p(\tilde{\bm{\xi}})-\frac{N}{|\mathbf{B}|}(\mathbf{\Theta}(\mathbf{B})^{\intercal}\,(\dot{\mathbf{b}}-\mathbf{\Theta}(\mathbf{B})\,\tilde{\bm{\xi}}$)), (7)
where $\dot{\mathbf{b}}\in\mathbb{R}^{|\mathbf{B}|}$ denotes the time derivative of a single dimension of the system states, corresponding to the column vector $\tilde{\bm{\xi}}\in\mathbb{R}^{m}$.
For an appropriately chosen step size schedule $\eta_{k}$, it has been established that the Langevin MCMC converges to the target stationary distribution [durmus2017nonasymptotic]. In practice, it may struggle to escape local traps when exploring multimodal distributions. To address these challenges, several methods, such as underdamped Langevin MCMC [cheng2018underdamped], Hamiltonian Monte Carlo [neal2011mcmc], cyclical SGMCMC [zhang2020cyclical], reSGLD [chen2018accelerating], etc., have been developed as extensions of the Langevin MCMC framework. These methods aim to improve computational efficiency and scalability, especially in high-dimensional settings. Related uncertainty-aware inference of Langevin dynamics has also been explored using Bayesian neural networks [bae2025inferring].
We consider reSGLD to balance exploration and exploitation when identifying system coefficients, which offers a blend of efficiency and robustness in traversing different energy levels to address large-scale non-convex sampling problems [deng2020non, zheng2024constrained]. It employs a high-temperature chain for exploration and a low-temperature chain for exploitation. The sampling process is shown as  [^2] Although reSGLD can admit multiple temperatures, we utilize two chains for simplicity.:
$\tilde{\bm{\xi}}_{k+1}^{(1)}$ $=\tilde{\bm{\xi}}_{k}^{(1)}-\eta_{k}\nabla\tilde{L}(\tilde{\bm{\xi}}_{k}^{(1)})+\sqrt{2\eta_{k}\tau_{1}}\bm{\zeta}_{k}^{(1)}$ (8)
$\tilde{\bm{\xi}}_{k+1}^{(2)}$ $=\tilde{\bm{\xi}}_{k}^{(2)}-\eta_{k}\nabla\tilde{L}(\tilde{\bm{\xi}}_{k}^{(2)})+\sqrt{2\eta_{k}\tau_{2}}\bm{\zeta}_{k}^{(2)},$
where $\tilde{\bm{\xi}}_{k}^{(1)}$ and $\tilde{\bm{\xi}}_{k}^{(2)}$ denote the sampling results of two chains with temperatures $\tau_{1}<\tau_{2}$. Furthermore, we swap the Markov chains in (8) with the corrected swapping function for the mini-batch setting, with the swap function defined as:
$\tilde{S}\big(\tilde{\bm{\xi}}_{k}^{(1)},\tilde{\bm{\xi}}_{k}^{(2)}\big)=e^{\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\left(\tilde{L}\big(\tilde{\bm{\xi}}_{k}^{(1)}\big)-\tilde{L}\big(\tilde{\bm{\xi}}_{k}^{(2)}\big)-\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\frac{\tilde{\sigma}^{2}}{C}\right)}$ (9)
where $\tilde{\sigma}^{2}$ approximates the variance of $\tilde{L}\big(\tilde{\bm{\xi}}_{k}^{(1)}\big)-\tilde{L}\big(\tilde{\bm{\xi}}_{k}^{(2)}\big)$ and $C$ acts as an adjustment to balance acceleration and bias. The algorithm proceeds iteratively until a specified number of iterations is reached. Generate uniform $u\in[0,1]$, if $u<\tilde{S}$, then perform the swapping between two chains. The parameter sets $\{\tilde{\bm{\xi}}_{k}^{(1)}\}_{k=1}^{K+1}$ are produced as outputs for analytical purposes.
While exact Bayesian inference can be computationally demanding, reSGLD leverages minibatch gradients and parallel chains to efficiently approximate the posterior distribution. This strategy avoids the restrictive assumptions of conjugate-prior Bayesian methods [fung2024] and provides richer uncertainty information, including multimodality and credible intervals.
3.2 Active Learning for Data-Efficient System Discovery
The integration of AL into data-driven discovery of dynamical systems aims to address two fundamental challenges in learning parsimonious dynamical models from limited and costly data. First, while accurate state derivatives $\mathbf{\dot{U}}$ are the cornerstone of the regression in Eq. (1), their acquisition is often a very expensive component of the discovery workflow. High-quality estimation requires dense, clean data, but collecting such data uniformly across the state space is highly inefficient and often infeasible. This highlights the critical need for a targeted data acquisition strategy with AL. Second, dynamical systems often exhibit heterogeneous information content across their state space. Sparse governing terms dominate in certain regimes (e.g., near bifurcations or transient states), while other regions contribute minimally to identifying the dynamical systems. Passive sampling strategies are inefficient and often fail to collect sufficient data in the information-rich regions. This oversight can lead to models that are not just inaccurate but fundamentally misrepresent the true system dynamics.
AL directly targets these issues by prioritizing data acquisition, where the uncertainty in the learned dynamics is maximized. By tackling (2) within a Bayesian framework, posterior uncertainties over the coefficients $\mathbf{\Xi}$ in (2) can be quantified. Regions of high predictive entropy or variance in $\mathbf{\dot{U}}$ correspond to states in which the model cannot confidently distinguish between candidate terms in the library $\mathbf{\Theta}(\mathbf{U})$. Actively querying derivatives in these states maximizes the information gain on $\mathbf{\Xi}$, which efficiently resolves ambiguities in the sparsity pattern [riis2023bayesian, pickering2022discovering, Fasel_2022]. This method, which acquires data with the highest predictive uncertainty, is also known as Active Learning MacKay (ALM) [MacKay1992, gramacy2009adaptive].
@@ 3.2 Active Learning for Data-Efficient System Discovery
To formulate AL, we first denote the pool of potential design points as $\mathcal{D}=\{\mathbf{U},\dot{\mathbf{U}}\}$, and denote the set of selected points as $\mathcal{M}$. With the information of the posterior $p(\mathbf{\Xi}|\mathbf{U})$, we sample $P$ groups of the model parameters from the posterior distribution and calculate $\dot{\mathbf{u}}^{pred}_{j}(t_{i})$ for each model $j$ at time $t_{i}$, with the mean of the prediction denoted as $\bar{\dot{\mathbf{u}}}^{pred}(t_{i})$. The predictive variance is calculated as
$\sigma^{2}({\dot{\mathbf{u}}}(t_{i}))=\frac{1}{P}\sum_{i=1}^{P}(\dot{\mathbf{u}}^{pred}_{j}(t_{i})-\bar{\dot{\mathbf{u}}}^{pred}(t_{i}))^{2},$
leading to the selection of potential design points with high uncertainty. In this paper, $\sigma^{2}(\cdot)$ is referred to as the uncertainty acquisition function. A key limitation observed in our experiments is that using the uncertainty acquisition function alone can lead to the redundant selection of points clustered in a single region. This occurs because while a query point may be informative, the uncertainty in its neighborhood often remains high, attracting subsequent samples. These clustered points offer diminishing returns in terms of new information about the overall system, resulting in a model that overfits the dynamics of one region while failing to learn from other potentially critical areas.
To address this clustering issue, BLADE adopts a hybrid acquisition strategy that balances uncertainty-driven exploitation with space-filling exploration. We augment the uncertainty measure with a density-adjusted maximin distance criterion [loeppky2010, yu2010]. This criterion promotes broader coverage by selecting new points that maximize the minimum distance to the existing dataset. The space-filling criterion is adjusted by the density of the data $\mathbf{u}(t_{i})$, and is defined as:
[…]
where the term $\lambda\geq 0$ is a hyperparameter that controls the strength of the density penalty applied to the distance measure. In the limiting case where $\lambda=0$, the penalty vanishes, and the criterion reverts to the original maximin design. A value of $\lambda=0.5$, for instance, balances the push for exploration (distance) with a preference for sampling in already populated regions (density). The objective of density adjustment is to penalize low-density points because they are more likely to have a large maximin distance, especially if they also have a large magnitude. Without this density adjustment, the search for space-filling points would be biased towards high-magnitude outliers and disregard potentially more informative points in dense clusters. This correction ensures the criterion promotes exploration across representative regions rather than simply selecting isolated points, and it also prevents overlap with the uncertainty-based component of our acquisition function.
Density-based adjustment has been adopted in AL research for both classification and regression tasks to achieve different goals, such as increasing the representativeness of the data [wang2021density], mitigating the sensitivity of the Euclidean distance to the scale and distribution of the data [donmez2007paired], and preventing outliers added to the training data [zhu2008active]. Several methods can be used to estimate data density, including Kernel Density Estimation (KDE), cosine distance-based metrics, and K-Nearest Neighbors (KNN). For this work, we select the KNN-based approach due to its advantages over the alternatives, which is calculated as the inverse of one point’s average distance to its K nearest neighbor. Compared to the computationally intensive KDE, KNN offers computational simplicity and directly captures local data density without evaluating kernel functions over all data. Furthermore, unlike cosine similarity, which captures only vector direction, KNN preserves both magnitude and directional information, which is crucial for identifying representative samples in our feature space.
The final acquisition function is constructed by combining the uncertainty and space-filling criteria. To ensure that they are on a comparable scale, both criteria are first standardized to the range $[0,1]$, yielding the normalized uncertainty $\tilde{\sigma}^{2}({\dot{\mathbf{u}}(t_{i})})$ and the normalized distance $\tilde{d}(\mathbf{\Theta}(\mathbf{u}(t_{i})))$. The final acquisition score is calculated as a weighted sum:
$C(\mathbf{u}(t_{i}))=\alpha\,\tilde{\sigma}^{2}({\dot{\mathbf{u}}}(t_{i}))+(1-\alpha)\,\tilde{d}(\mathbf{\Theta}(\mathbf{u}(t_{i}))),$ (11)
where a tunable hyperparameter $\alpha\in[0,1]$ balances their relative contributions to the final acquisition score.
@@
Input: Pool of potential design points $\mathcal{D}$; number of initial points $n$; batch size of selected points per round $m$, maximum allowed sample size $N_{max}$, tolerance of convergence Tol.
@@
Randomly select $n$ points, $\mathbf{u}(t_{1}),\dots,\mathbf{u}(t_{n})\in\mathbf{U}$ with derivatives $\dot{\mathbf{u}}(t_{1}),\dots,\dot{\mathbf{u}}(t_{n})$ available;
Set of selected points $\mathcal{M}=\{(\mathbf{u}(t_{1}),\dot{\mathbf{u}}(t_{1})),\dots,(\mathbf{u}(t_{n}),\dot{\mathbf{u}}(t_{n}))\}$, $\mathcal{D}\leftarrow\mathcal{D}/\mathcal{M}$ // i.e., exclude selected data from the design pool;
@@
   Use $\mathcal{M}$ points to construct the library $\mathbf{\Theta}$ and perform BLADE to estimate the posterior distribution $p(\bm{\mathbf{\Xi}}\mid\mathcal{M})$;
@@
Algorithm 1 Active Learning in BLADE: Iteratively improve the model by selecting design points that maximize the acquisition function, updating the posterior, and retraining the model until convergence.
3.3 The Proposed Algorithm